# Identity-Obscured Neural Radiance Fields: Privacy-Preserving 3D Facial Reconstruction

### v1

@@ 3.2 Two-stage Training
In practical implementation, this operator can take the form of an edge extraction operator, such as the Sobel operator. Its design aims for versatility, enabling users to choose from a range of operators $\sigma$ depending on their specific privacy requirements. This flexibility empowers users to acquire data with different levels of privacy protection.
During this stage, as there is no RGB information for supervision, MLP predictions for color values may exhibit errors. This is another aspect affirming the privacy-preserving nature of our method. At the same time, since we introduce alternative forms of supervision for color variations, facial geometric details can be recovered even when color information in the rendered images seems unreasonable. Our primary focus is on the recovery of geometric information, with the goal of generating a reasonable geometry for downstream tasks. In cases where the camera viewpoint is in front of the face, the training process solely involves the supervision of gradient components to ensure privacy preservation.
3.3 Privacy-preserving argumentations
@@
The purpose of the preserving operator is to truncate the user’s private information directly on the user’s end, without exposing it to any server. This stage requires both simplicity and convenience, while ensuring the highest level of privacy protection. To safeguard user privacy, we conduct an analysis in three key aspects using the preserving operator:
@@
To safeguard privacy, we retain only color variation details and discard complete RGB data from color images. This privacy-conscious approach preserves essential data for reconstruction. During image uploads, we retain only the gradient magnitude information $\left\|\mathbf{g}\right\|$, making it practically impossible to reverse engineer RGB data due to many-to-one mapping relationships.
@@
We provide an example in Figure 4, illustrating that multiple images may correlate with the same gradient amplitude information. In other words, a single color gradient size can correspond to numerous color gradient values and color values. This scenario implies that multiple images can share the same gradient magnitude information, making the retrieval of the original pixel values a formidable challenge. This lack of one-to-one correspondence adds a layer of complexity to privacy breach attempts.
Figure 4: Due to the nature of vector magnitude, the blurry color gradient modulus derived from the low-resolution GT image can correspond to numerous other images, including those on the left, and not solely the GT image.
@@
After gradient processing, the data becomes visually indistinguishable to human eyes. This transformation not only enhances privacy but also results in highly similar images for different identities. Learned Perceptual Image Patch Similarity (LPIPS) is a perceptual similarity measure based on learning, which is more in line with human perception. We use LPIPS to measure the similarity of images before and after privacy protection, as quantified in Table 1. This processing ensures that facial images are privacy-preserving while maintaining their availability for server-side geometric reconstruction tasks.
@@
Table 1: LPIPS from different identities shows that the preserving images are much more similar.
3.4 Optimization
Template.
Templates offer an effective solution when users cannot capture photos without privacy-sensitive information, such as handheld devices or other constraints. In such cases, user-uploaded images need processing for privacy protection. To address this challenge, we propose using a non-sensitive head template. This enables users to recover head contour geometry even when privacy-neutral images are scarce, facilitating effective geometric reconstruction. While template usage is not mandatory, it enhances our model’s compatibility with various data inputs, making it more versatile for a smoother user experience.
Regularization.
In our network structure, based on the neural radiance fields pipeline, we make specific optimizations to enhance geometric representation. Our primary goal is not the final rendering appearance within the pipeline but the recovery of fine geometric details for downstream tasks. To achieve this, we apply regularization constraints to the color rendering network. These constraints encourage the network to prioritize learning geometric intricacies. In previous work [26], a color regularization constraint was proposed which introduced trainable bounds for each layer, effectively constraining the expression of MLP layers using knowledge from Lipschitz continuous networks. In the specific network structure, each MLP layer is reformulated as $y=\sigma(\widehat{W}_{i}\mathbf{x}+b_{i})$, and $\widehat{W}_{i}=a(W_{i},\text{softplus}(k_{i}))$. During the privacy-preserving stage, we apply color regularization by constraining the product of Lipschitz constants, $k_{i}$, for each layer:
$\text{softplus}(k_{i})=\ln{(1+e^{k_{i}})}.$ (7)
$\mathcal{L}_{\text{lip}}={\textstyle\prod_{i=1}^{l}}\text{softplus}(k_{i}).$ (8)
Putting it all together, our training loss is as follows:
ID 340 ID 421 ID 393
Refer to caption Refer to caption Refer to caption
Figure 5: Due to the absence of color supervision in Stage 2, only unrealistic RGB images can be rendered. The impractical nature of these images emphasizes the underlying motivation for our work in safeguarding privacy.
4 Experiments
4.1 Setup
Datasets.
In our experiments, we employ two separate datasets, with each identity’s data consisting of either 30 or 36 RGB images with a resolution of $64\times 64$ pixels. The decision to opt for a lower resolution and increased blurring is intended to augment privacy protection. We utilize the FaceScape [38] and the High-Fidelity 3D Head (H3DS) [25] datasets in our experiments. FaceScape, a novel method for rendering textured 3D faces, provides data for various subjects and their specific expressions. The H3DS dataset [25] includes headshot data from various countries, encompassing diverse hair and skin types. We randomly sampled 10 identities from each of the two datasets for our experiments. Both datasets provide essential information on camera poses and associated parameters, with certain datasets providing masks to facilitate simplified data extraction. the selection of images for training should be adaptable, taking into account the user’s privacy preferences and concerns. Here, we present an example as illustrated in the above inset. In our subsequent experiments with the FaceScape dataset, our exclusive focus is on a fixed set of 10 privacy-neutral views, while the remaining 20 views are considered as privacy-preserving images. In the H3DS dataset, we also adopted the selection criteria mentioned in the example to choose 16 privacy-neutral images and 20 images necessitating privacy preservation. Through experiments conducted on these datasets, we demonstrate the adaptability and robustness of our method in reconstructing facial data across various scenarios.
Baseline.
Our approach is based on VolSDF [41], and we evaluate the geometric accuracy of our work under identical camera poses and view directions. We have also provided the rendered results in Figure 5, confirming that the generated images are not realistic and do not compromise people’s privacy. VolSDF is based on all full-face images, while our method exclusively utilizes privacy-preserving data, as illustrated in the example shown in the above inset. Given that our method is trained on gradient images, a large amount of information is lost compared to full RGB images at all viewing angles, and our method still achieves results comparable to VolSDF, which will be specified in the Evaluation.
GT image Stage 1 Stage 2 VolSDF (full img) GT mesh
ID 346 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
ID 04 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: We propose a two-stage strategy. The first stage focuses on learning the basic geometry of the head, and the second stage only uses privacy-preserving data to optimize facial geometric details. Even compared with the full RGB image given 30 viewing directions, our reconstruction still achieves comparable results.
Evaluation.
To ensure fair experiments, we maintain the same training epochs (1.5k) and employ $512\times 512$ resolution matching cubes to extract the final mesh. The mesh evaluation is conducted by appropriately cropping it to the size of the provided GT mesh, ensuring a fair assessment of facial geometric reconstruction accuracy. We evaluate the quality of our algorithm’s geometry recovery through both qualitative and quantitative comparisons with the VolSDF algorithm. We present results obtained after the first stage of privacy-neutral training and the second stage of privacy-preserving reconstruction, comparing them with VolSDF’s full-image supervised reconstruction outcomes in Figure 6. To evaluate the overall reconstruction accuracy, we employ the Chamfer distance (CD) metric, computed for both our method and VolSDF. These metrics are detailed in Table 2. Furthermore, we extend our comparisons to include two additional datasets with braoder viewing angles, extending beyond facial features. In these cases, our reconstruction excels in capturing facial details such as hair in Figure 7 or interesting facial features such as different expressions in Figure 8. This highlights the versatility of our method in handling intricate aspects beyond facial contours.
[…]
VolSDF [41](full img) $0.4509$ $0.5163$ $0.5399$ $0.4547$ $0.4988$ $0.5266$ $0.5059$ $0.4366$ $0.5394$ $0.4788$ $0.4948$
[…]
(b) CD (mm) results on 10 identities in the H3DS dataset [25].
ID 08 ID 00 ID 03
Refer to caption Refer to caption Refer to caption
Figure 7: In terms of reconstruction quality, even with 64-resolution images, we are still able to capture many geometric details, including hair.
@@ 4.1 Setup
4.2 Ablation studies
Lipschitz regularization.
We conduct an ablation study of the Lipschitz regularization proposed in our paper, as depicted in Figure 9. Although a face represents a relatively smooth geometry, it still possesses intricate geometric details, such as wrinkles. The Lipschitz regularization plays a crucial role in enhancing these facial details, ultimately resulting in the lowest CD. We apply this regularization in the second stage to focus on refining the geometric intricacies of the faces.
ground truth w/o Lip. w/ Lip.
ID 393 Refer to caption Refer to caption Refer to caption
ID 421 Refer to caption Refer to caption Refer to caption
ID 395 Refer to caption Refer to caption Refer to caption
CD ($10^{-3}$) ID 393 ID 421 ID 395 w/ Lip. 0.6998 0.7176 0.7664 w/o Lip. 0.7753 0.7864 0.8347
Figure 9: Lipschitz regularization (Lip.) makes geometry learning more reasonable in Stage 2, allowing the smoothness of the face to be maintained while retaining clear features and details, such as the mouth (Row 1) or the bridge of the nose (Row 3).
Sparse view with template.
Considering that users may have varying definitions of privacy data, we also provide a comparison of scenarios with fewer color images in the first stage, as illustrated in Figure 10. It is evident that when users have stricter privacy definitions during the first stage, resulting in limited data acquisition, it may impact the reconstruction accuracy of the results of stage 1. Nevertheless, the impact on the final result remains acceptable.
In extreme cases, where users do not provide privacy-neutral images, we can still perform a one-stage geometric simulation using the privacy-neutral multi-head template. The results are shown in Figure 16. The provision of identity-free geometry templates cannot completely replace the original reconstruction, because the missing and inconsistent contours will increase the difficulty of subsequent training to a certain extent. Nevertheless, it proves highly effective in addressing special situations.
Stage 1 Stage 2
Figure 10: The input of varying views and different quantities of RGB images in the first stage may influence the reconstruction accuracy of Stage 1, but it has minimal impact on the final reconstruction quality. Row 1 demonstrates a one-stage process under a 15-view setting, while Row 2 and Row 3, both under a 10-view setting, exhibit different input views.
Template Results
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Template. Two on the left showcase the neutral template we employ, acquired through training with multiple heads. two on the right are our reconstruction results using the template. This highlights that privacy-neutral templates contribute to reconstruction without the need of privacy-neutral images.
Critical role of initial geometric priors.
In our experiments, the first stage plays a critical role, as it involves extracting essential geometric prior knowledge about the human head. The accuracy of the initial outline, closely resembling the person, significantly benefits the second stage of the process. We also provide results for scenarios without the first stage. When using all the gradient reconstruction examples, we observe that the reconstruction lacks smoothness and does not represent a plausible head geometry in Figure 12 (a). This is due to the absence of low-frequency geometric information at the foundation. Moreover, if we jointly train privacy-neutral and privacy-preserving data on a template, as depicted in Figure 12 (b), the results remain unsatisfactory despite training the model for the same number of epochs in the two-stage process. This is because the geometric low-frequency information has not been adequately established during the initial training, and the introduction of high-frequency data disrupts the overall learning process of the model.
(a) 27.65 (b) 15.67 (c) 0.84
Figure 12: Training using both rgb and gradient losses in a single stage fails to achieve satisfactory reconstruction, as evidenced by the high CD ($10^{-3}$) values. (a) shows the single-stage reconstruction using only gradient loss; (b) shows the single-stage training with both rgb and gradient losses; (c) shows our two-stage training result.
5 Conclusion
In this paper, we emphasize the significance of privacy in neural facial reconstruction, an aspect that is often overlooked. We introduce a method capable of reconstructing detailed head geometry, particularly focusing on facial features, from identity-obscured input. Our approach involves classifying facial data, applying privacy-destroying processing to images containing sensitive facial information, retaining only the parts crucial for geometric reconstruction, and processing the two types of data separately in a two-stage manner. Extensive experiments demonstrate the effective balance we achieve between privacy protection goals and reconstruction accuracy. Our methodology marks a significant advancement by seamlessly integrating the realms of privacy protection and neural facial reconstruction, ushering in a new era of exploration in this domain.
Supplementary Material
6 Discussions
Privacy Risk Assessment.
Our overarching objective is the utilization of identity-obscured input for reconstruction tasks, ensuring the imperviousness of private information as images traverse beyond the user terminal. This involves an initial down-sampling of all images, reducing them to a 64-resolution level. Subsequent to this, a privacy-preserving operator is applied to transform these images into privacy-neutral counterparts, as illustrated in Figure 13. Importantly, this process effectively precludes the extraction of private information from the images.
In exploring information related to image privacy, we referred to a set of attributes proposed by [22]. This research is grounded in substantial data on user privacy preferences and privacy ratings, culminating in a predictive model that infers privacy attributes and specific risks associated with these attributes. The model is calibrated with 68 categories, encompassing attributes such as credit card, tattoo, eye color, age group, and more. Utilizing this standardized model, we evaluate privacy risks, as illustrated in Figure 13. Following the privacy protection methods we proposed, these privacy attributes in the raw image become unidentifiable. This emphasizes the inherent value of the problem we have addressed. Due to the existence of multiple classes for privacy attributes, privacy leakage is assessed using the mean average precision averaged across these classes. This metric is also employed to evaluate privacy preservation within the framework of a learning anonymization function, as highlighted in studies such as [6, 8]. These studies rely on this algorithm to compute privacy leakage, thereby evaluating the privacy risk associated with videos after privacy protection. Under the premise of privacy protection, our reconstruction task continues to achieve results comparable to existing methods that reconstruct full images.
Figure 13: Single image of identity 340 showing the type of private attributes (shown in red on the left image) leaked in visual data. The right panel displays an image with obscured identity, where these attributes are scarcely discernible. Remarkably, our method maintains a parallel level of reconstruction performance while significantly mitigating the visibility of private attributes.
7 Implementations
Datasets.

### v2

@@ Introduction
In practical implementation, this operator can take the form of an edge extraction operator, such as the Sobel operator. Its design aims for versatility, enabling users to choose from a range of operators $\sigma$ depending on their specific privacy requirements and to adjust operator parameters to ensure the irreproducibility of the reconstruction. This flexibility empowers users to acquire data with different levels of protection. During this stage, the absence of RGB supervision may lead to inaccuracies in MLP color predictions. Nevertheless, our method effectively recovers facial geometric details even when radiance information appears unreliable.
The privacy-preserving approach focuses on three key aspects: irreversibility by retaining only gradient magnitudes, color multiplicity where different images can map to the same gradient, and perceptual indistinction to keep the processed data visually indistinguishable from the original. Detailed discussions are in the supplementary material.
Optimization
Template.
Templates offer an effective solution when users cannot capture photos without privacy-sensitive information, such as handheld devices or other constraints. In such cases, user-uploaded images need processing for privacy protection. To address this challenge, we propose using a neutral head template. This enables users to recover head geometry even when privacy-neutral images are scarce, facilitating effective geometric reconstruction. While template usage is not mandatory, it enhances our model’s compatibility with various data inputs, making it more versatile for a smoother user experience.
Regularization.
In our network structure, based on the neural radiance fields pipeline, we make specific optimizations to enhance geometric representation. Our primary goal is not the final rendering appearance but the recovery of fine geometric details for downstream tasks. To achieve this, we apply regularization constraints to the color rendering network. These constraints encourage the network to prioritize learning geometric intricacies. In previous work (Rosu and Behnke 2023), a color regularization constraint was proposed which introduced trainable bounds for each layer, effectively constraining the expression of MLP layers using knowledge from Lipschitz continuous networks. In the specific network structure, each MLP layer is reformulated as $y=\sigma(\widehat{W}_{i}\mathbf{x}+b_{i})$, and $\widehat{W}_{i}=a(W_{i},\text{softplus}(m_{i}))$, where $a$ normalizes the weight matrix. During the privacy-protected stage, we apply color regularization by constraining the product of Lipschitz constants, $m_{i}$, for each layer:
$\text{softplus}(m_{i})=\ln{(1+e^{m_{i}})}.$ (7)
$\mathcal{L}_{\text{lip}}={\textstyle\prod_{i=1}^{l}}\text{softplus}(m_{i}).$ (8)
Putting it all together, our training loss is as follows:
Experiments
Setup
Datasets.
In our experiments, we utilize two representative datasets: FaceScape (Yang et al. 2020) and High-Fidelity 3D Head (H3DS) (Ramon et al. 2021). Each dataset includes 30 to 36 RGB images per identity at $64\times 64$ pixels, with lower resolution and increased blurring chosen to enhance privacy protection. FaceScape offers textured 3D face data for various subjects and expressions, while H3DS, captured in real-world scenarios, provides headshot data across different countries, ethnic backgrounds, and lighting conditions. We randomly sampled 10 identities from each dataset and treated all images with visible facial features as sensitive images. The selection of images for training is adaptable to user privacy preferences, as detailed in the supplementary material. In our experiments with the FaceScape dataset, we utilize 10 images that are inherently privacy-neutral and process 20 images as privacy-protected. Similarly, for the H3DS dataset, 16 images are inherently privacy-neutral, and 20 images are processed as privacy-protected. While our primary analysis focuses on these two datasets, we also test additional facial datasets to broaden our validation in the supplementary. These experiments demonstrate the adaptability and robustness of our method in various conditions.
GT image Stage 1 Stage 2 VolSDF (full img) GT mesh
Figure 2: A two-stage strategy: stage 1 focuses on learning the basic geometry of the head, and stage 2 only uses gradient images to optimize facial geometric details. Even compared with the full RGB image given 30 viewing directions, our reconstruction still achieves comparable results.
Baseline.
Our approach is based on VolSDF (Yariv et al. 2021) due to its stability in constructing detailed facial geometry. The flexibility of our framework allows for the incorporation of alternative rendering and training methods such as 2DGS (Huang et al. 2024) in Figure 11. We evaluate the geometric accuracy of our work under identical camera poses and view directions. VolSDF is based on all full-face images, while our method doesn’t rely on sensitive facial images. Given that our method is trained on gradient images, a large amount of information is lost compared to full RGB images at all viewing angles, and our results remain comparable to VolSDF, as detailed in the Evaluation.
Evaluation.
We evaluate the quality of our algorithm’s geometry recovery through both qualitative and quantitative comparisons with the VolSDF algorithm. We present results obtained after the first stage of privacy-neutral training and the second stage of privacy-protected reconstruction, comparing them with VolSDF’s full-image supervised reconstruction outcomes in Figure 2. To evaluate the overall reconstruction accuracy, we employ the Chamfer distance (CD) metric, computed for both our method and VolSDF. These metrics are detailed in Table 1. Furthermore, our reconstruction excels in capturing intricate details such as hair or interesting facial features such as different expressions in Figure 3. This highlights the versatility of our method in handling intricate aspects beyond facial contours.
[…]
VolSDF (Yariv et al. 2021)(full img) $0.4509$ $0.5163$ $0.5399$ $0.4547$ $0.4988$ $0.5266$ $0.5059$ $0.4366$ $0.5394$ $0.4788$ $0.4948$
[…]
(b) CD (mm) results on 10 identities in the H3DS dataset (Ramon et al. 2021).
ID 08 ID 00 ID 03
Refer to caption Refer to caption Refer to caption
  angry shock smile
Refer to caption Refer to caption Refer to caption
Figure 3: In terms of reconstruction quality, even with 64-resolution images, we are still able to capture many geometric details, including hair and expressions.
ground truth w/o Lip. w/ Lip.
ID 393 Refer to caption Refer to caption Refer to caption
ID 395 Refer to caption Refer to caption Refer to caption
CD ($10^{-3}$) ID 393 ID 395 w/ Lip. 0.6998 0.7664 w/o Lip. 0.7753 0.8347
Figure 4: Lipschitz regularization (Lip.) makes geometry learning more reasonable in Stage 2, such as the mouth (Row 1) and the smoother face (Row 2).
Ablation studies
Lipschitz regularization.
We conduct an ablation study of the Lipschitz regularization as depicted in Figure 4. Since the face represents relatively smooth geometry, Lipschitz regularization helps ensure more accurate geometric learning and plays a crucial role in recovering facial details, ultimately resulting in the lowest CD.
Sparse views and geometric priors.
Considering varying user definitions of privacy, we compare scenarios with fewer color images in Stage 1, as shown in Figure 5. Stricter privacy requirements may impact Stage 1 accuracy, but the final results remain acceptable. In extreme cases where all provided images contain sensitive features, a one-stage geometric simulation using identity-free templates trained on multiple heads can be used. Figure 6 shows that these templates, while not fully replacing original reconstructions due to missing low-frequency geometric information, are effective in special situations.
Our experiments show that the first stage is crucial for extracting essential geometric priors, benefiting the second stage significantly. Without this initial stage, reconstructions from gradient examples lack smoothness and plausible geometry, as shown in Figure 7 (a), due to missing low-frequency information. Training with both privacy-neutral and privacy-protected images on a template, as shown in Figure 7 (b), yields unsatisfactory results. This occurs because inadequate low-frequency information during the initial training disrupts the learning process when high-frequency data is introduced.
Stage 1 Stage 2
Figure 5: Varying the number of RGB images and views in the first stage may affect Stage 1 accuracy but has minimal impact on final reconstruction quality. Row 1 shows a one-stage process with 15 views, while Rows 2 and 3, both with 10 views, display different input views.
Template Results
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Template. Two on the left showcase the neutral template we employ, while two on the right are our reconstruction results using the template.
(a) 27.65 (b) 15.67 (c) 0.84
Figure 7: Single stage reconstruction with CD ($10^{-3}$) values. (a) shows the single-stage reconstruction using only gradient loss; (b) shows the single-stage training with both rgb and gradient losses; (c) shows our two-stage training result.
Security Discussion
Avoiding 2D FR tracking.
Method Original Protected Rendered Mesh Our Rendering VolSDF Rendering
VGG-Face 93.75 6.51 3.03 43.74 84.62
Facenet 96.88 9.68 18.18 31.23 81.52
Facenet512 96.87 3.23 3.20 18.77 79.27
ArcFace 90.63 6.66 6.46 31.25 79.23
Dlib 93.75 3.17 1.77 37.50 71.57
SFace 84.38 6.45 9.14 43.71 73.35
GhostFaceNet 85.42 6.62 1.04 31.44 73.92
Average 91.67 6.05 6.12 33.95 77.64
Table 2: Comparison of recognition accuracy (%) across 7 FR systems shows that our protected images, directly rendered mesh images, and our rendered results all significantly reduce recognition accuracy, thus avoiding FR tracking.
Although 3D meshes contain identifiable information, they are more robust against malicious exploitation compared to 2D facial images. This robustness stems from the challenges in collecting large-scale 3D datasets, which has limited the development of deep learning-based 3D face recognition (Guo et al. 2023). In contrast, most widely deployed FR systems rely on 2D images due to their speed and convenience. The powerful capabilities of 2D FR systems, if misused for tracking and surveillance, can lead to severe privacy concerns. To demonstrate the effectiveness of our privacy-preserving method, we evaluated its resistance against seven state-of-the-art (SOTA) 2D face recognition systems (Parkhi, Vedaldi, and Zisserman 2015; Schroff, Kalenichenko, and Philbin 2015; Deng et al. 2019; King 2009; Zhong et al. 2021; Alansari et al. 2023), which typically represent faces as vectors, using metrics like cosine similarity (Figure 8) to measure the similarity between images of the same person. The tests results in Table 2 included original sensitive images (Original), protected images (Protected), rendering results from both our method (Our Rendering) and VolSDF (VolSDF Rendering), and images directly rendered from 3D geometric meshes (Rendered Mesh). which confirms that our approach significantly reduces FR accuracy, demonstrating its effectiveness in preventing tracking and lowering security risks.
Original 11 Original 11 Original 08 Protected 11 Mesh 11
Verified true false false false Distance 0.4978 0.6145 0.9095 0.8893
Figure 8: VGG verification of the four images (right) against the original (left), showing cosine distances. Greater distances indicate lower similarity.
GT VolSDF Ours Refer to caption
Source Target Origin VolSDF Ours Refer to caption
Figure 9: We reconstruct geometric details without rendering realistic sensitive images (row 1). Our rendered images after DeepFake processing retain significantly different facial details from the target person (row 2,3), demonstrating that DeepFake fails to convincingly swap faces with our results thus our method effectively prevents impersonation.
Counteracting DeepFake.
Current DeepFake technology can convincingly replace facial details through other people’s images and videos, resulting in highly realistic and potentially harmful fabrications. In NeRF-based reconstructions, DeepFake can participate in fabrications in two ways: one requires users to provide multi-view facial images, which our method avoids, and the other applies DeepFake to rendered images. Figure 9 compares results from DeepFake models with (Row 2) and without (Row 3) pre-trained generative models. The inputs include original multi-view RGB images, images rendered with VolSDF, and images rendered with our method. Both types of our DeepFake images fail to convincingly replicate the target individual due to the loss of sensitive facial textures, with quantitative analysis provided in the supplementary materials, making DeepFake technology ineffective.
Preventing misuse.
We investigated the role of facial texture in visual recognition by transferring the radiance and geometry from different identities. Figure 10 demonstrated that facial mesh transferred with someone else’s facial texture radiance appeared significantly more similar to that person, reinforcing the idea that achieving accurate forgery of facial identities requires the acquisition of highly specific, original facial textures. Our findings highlight that without access to the exact texture of the original face, the potential for successful face forgery remains limited.
Original Geometry Transfer Results Transfer
[…]
Future application.
Full images reconstruction (2DGS) Protected images reconstruction (2DGS)
Figure 11: For sensitive images (license plates, conference room details), users can provide only protected inputs instead of the actual images, and reconstruct details.
For 3D reconstruction tasks, our proposed method prevents reliance on sensitive images by achieving high-fidelity geometric reconstruction using only sensitive-image gradients as inputs. This finding is applicable to other neural reconstruction pipelines (Huang et al. 2024) and extends beyond facial reconstruction to various scenarios requiring minimal input such as meeting rooms or scenes involving protected elements like vehicle license plates (Figure 11). Notably, even with such minimal input, the quality of geometric reconstruction remains comparable.
Conclusion
In this paper, we highlight the often-overlooked aspect of privacy in neural facial reconstruction. We present a method that reconstructs detailed head geometry without relying on sensitive input. By processing sensitive facial data and focusing on essential geometric information in a two-stage process, our approach balances privacy protection with reconstruction quality. Our method represents a significant advancement, integrating privacy and neural facial reconstruction and paving the way for new explorations in this field.
Appendix
Privacy Quantitative experiments
We have provided visual results in the paper and will discuss the findings in detail, along with presenting quantitative results to further demonstrate the effectiveness of our privacy-protection approach.
DeepFake
Deep learning has revolutionized computer vision, enabling highly realistic digital image manipulation. A prominent example is face swapping, where a source face is seamlessly transferred onto a target, retaining the target’s facial movements and expressions. DeepFake technology, now highly advanced, can achieve film-quality effects, with DeepFakeLab (Perov et al. 2020) being the leading software for creating such content. In our study, we first trained the model for 80k iterations without the assistance of a generative model to swap Elon Musk’s face. Second, with the rise of generative models, many DeepFake techniques now utilize source expressions and attributes while applying new facial features to make the generated face closely resemble the target individual, so we also conducted face swaps using a generative model. In our paper, we present two sets of results, comparing cases with and without the generative model. Visually, the high realism of neural rendering allows both VolSDF-rendered outputs and the original input images to be clearly identified as the target identity in DeepFakes. However, our rendered results do not contain sensitive facial features, and the DeepFakes generated from them lack clear identity-specific traits.
Face recognition systems (Parkhi, Vedaldi, and Zisserman 2015; Schroff, Kalenichenko, and Philbin 2015; Amos et al. 2016; King 2009; Zhong et al. 2021; Alansari et al. 2023) can measure similarity using cosine distance, where a smaller value indicates a higher similarity between two images. For specific implementation details, we referred to (Serengil and Ozpinar 2020). For more results and quantitative analysis, please refer to Table 3. DeepFake videos can be found in the supplementary video we provide.
Source Target Origin VolSDF Ours
Method Original VolSDF Our VGG-Face 0.5697 0.5758 0.7431 Facenet 0.6132 0.5794 0.659 Facenet512 0.4068 0.4345 0.5973 OpenFace 0.3548 0.3566 0.4112 ArcFace 0.3523 0.3918 0.9121 Dlib 0.0413 0.0391 0.1153 SFace 0.7191 0.7154 0.7513 GhostFaceNet 0.6288 0.6618 0.9541
(a) DeepFake without generation model.
Source Target Origin VolSDF Ours
Method Original VolSDF Our VGG-Face 0.5881 0.6133 0.668 Facenet 0.6601 0.6365 0.4443 Facenet512 0.3294 0.3351 0.4744 OpenFace 0.3732 0.3482 0.4237 ArcFace 0.3345 0.3166 0.4542 Dlib 0.0478 0.0422 0.0661 SFace 0.5457 0.4976 0.5953 GhostFaceNet 0.3538 0.3665 0.5030
(b) DeepFake with generation model.
Source Target Origin VolSDF Ours
Method Original VolSDF Our VGG-Face 0.468 0.6018 0.6027 Facenet 0.2067 0.3735 0.5196 Facenet512 0.1376 0.3197 0.4668 OpenFace 0.1632 0.2264 0.3766 ArcFace 0.3247 0.5336 0.667 Dlib 0.0502 0.0781 0.0695 SFace 0.5342 0.6945 0.6786 GhostFaceNet 0.3895 0.5549 0.6571
(c) DeepFake without generation model.
Source Target Origin VolSDF Ours
Method Original VolSDF Our VGG-Face 0.6913 0.7852 0.8501 Facenet 0.359 0.5103 0.4897 Facenet512 0.3467 0.5215 0.536 OpenFace 0.1966 0.3026 0.352 ArcFace 0.4838 0.5907 0.6794 Dlib 0.0435 0.051 0.0867 SFace 0.6056 0.6127 0.7718 GhostFaceNet 0.4732 0.6102 0.7356
(d) DeepFake with generation model.
Table 3: Comparison of the original image, original image VolSDF rendering, and our rendering after DeepFake, with both quantitative and qualitative analysis. A higher value indicates a lower similarity to the target. Our method often achieves the lowest similarity, effectively preventing DeepFake.
Appearance Transfer
Original State:
Person A has facial geometry $\text{Person A has facial geometry }G_{A}\text{ and radiance information }R_{A}.$
Person B has facial geometry $\text{Person B has facial geometry }G_{B}\text{ and radiance information }R_{B}.$
After Appearance Transfer:
$A^{\prime}=\text{Transfer}(G_{A},R_{B})$
$B^{\prime}=\text{Transfer}(G_{B},R_{A})$
Visual Changes:
[…]
We employ the method proposed by (Xu et al. 2023) for achieving appearance transfer. This approach is straightforward and does not require additional operations; it simply aligns the radiance and geometry information in three dimensions without undergoing any additional retraining, ensuring the fairness of this alignment.
In the Preventing Misuse section, we discussed that without accurate sensitive facial features, even obtaining the geometry alone does not enable forgery. Here, we provide specific cosine distance values in Figure 12 to demonstrate that, in addition to visual similarity, the transferred results are numerically closer to the radiance results. This further confirms that our proposed method effectively prevents misuse by avoiding the inclusion of sensitive facial features.
Transferred Image Geometry provider Radiance provider
G401&R429 401 429
Refer to caption Refer to caption Refer to caption
False / 0.6360 True / 0.2570 G401&R491 401 491
Refer to caption Refer to caption Refer to caption
False / 0.4295 True / 0.1109 G429&R401 429 401
Refer to caption Refer to caption Refer to caption
False / 0.4728 True / 0.2308 G429&R491 429 491
Refer to caption Refer to caption Refer to caption
False / 0.5468 True / 0.2183 G491&R429 491 429
Refer to caption Refer to caption Refer to caption
False / 0.5964 True / 0.1534 G491&R401 491 401
Refer to caption Refer to caption Refer to caption
False / 0.4804 True / 0.2512
Figure 12: The values represent the distances to the transferred image, a smaller distance means more similarity. After transferring, the face is recognized as the identity of the skin texture rather than its original geometry provider, as evaluated using Facenet (Schroff, Kalenichenko, and Philbin 2015).
Protecting Operators
Privacy-protected argumentations
@@ Conclusion
The purpose of the preserving operator is to truncate the user’s private information directly on the user’s end, without exposing it to anyone. This process requires both simplicity and convenience while ensuring the highest level of privacy protection. To safeguard user privacy, we analyze three key aspects using the preserving operator:
@@
To safeguard privacy, we retain only color variation and discard complete RGB data from color images. This privacy-conscious approach preserves essential data for reconstruction. During image uploads, we retain only the gradient magnitude information $\left\|\mathbf{g}\right\|$, making it practically impossible to reverse engineer RGB data due to many-to-one mapping relationships.
@@
We provide an example in Figure 14, illustrating that multiple images may correlate with the same gradient amplitude information. In other words, a single color gradient module can correspond to numerous color gradient values and color values. This scenario implies that multiple images can share the same gradient magnitude information, making the retrieval of the original pixel values a formidable challenge. This lack of one-to-one correspondence adds a layer of complexity to privacy breach attempts.
Figure 14: The blurry color gradient modulus derived from the low-resolution GT image can correspond to numerous other images, including those on the left, and not solely the GT image.
@@
After gradient processing, the data becomes visually indistinguishable to human eyes. This transformation not only enhances privacy but also results in highly similar images for different identities. Learned Perceptual Image Patch Similarity (LPIPS) is a measure of perceptual similarity based on learning that is more in line with human perception. We use LPIPS to measure the similarity of images before and after privacy protection, as quantified in Table 4. This processing ensures that facial images are privacy-protected while maintaining their availability for geometric reconstruction tasks.
@@
Table 4: LPIPS from different identities shows that the privacy-protected images are much more similar.
A broader evaluation
Robustness
These datasets we used in the paper cover a wide range of information, encompassing distinct lighting conditions (shown in Figure 15) and individuals of various races, skin tones, and ages. This diversity highlights the success of our experiments in achieving detailed geometric reconstruction across a spectrum of conditions.
Figure 15: Our experiments are evaluated under different lighting conditions, especially the H3DS dataset.
We have introduced a highly flexible framework that currently utilizes a two-stage pipeline based on based on VolSDF (Yariv et al. 2021) for reconstruction. To address additional geometric privacy scenarios, such as private meeting rooms or license plate reconstruction mentioned in our paper, we have adopted other neural reconstruction methods, such as 2DGS (Huang et al. 2024), that are capable of managing these 3D scenes. Therefore, we encourage the exploration of neural rendering accelerators proposed in current research, such as (Müller et al. 2022; Huang et al. 2024), for potential optimizations in future work.
Expressions
Our model demonstrates robust performance across a wide range of facial expressions, showcasing its ability to effectively capture and represent subtle emotional nuances 16. This versatility highlights the model’s potential for diverse applications, including advanced emotion recognition, and interactive AI systems. The expanded results underscore that our model is not only adept at handling common facial expressions but also offers detailed and nuanced representations, such as emotional states, which could be leveraged in innovative ways across various domains.
@@
Implementation
Datasets